please help me to fin the answer of this question : ab+a-b=2001, find the value of a and b. if a and b are a real number.
Answer by Alan3354(69443)

please help me to fin the answer of this question : ab+a-b=2001, find the value of a and b. if a and b are a real number.
---------------------
You have 1 equation in 2 unknowns, so no solution is possible. You can only find one variable in terms of the other:
ab+a-b=2001
a(*b+1)-b = 2001
a = (2001+b)/(b+1)
Or,
b*(a-1)+a=2001
b = (2001-a)/(a-1)